The weights (lb) of discarded plastic from a random sample of 62 household has sample mean of


"\bar{x}=1.911"


and a sample standard deviation of

"s=1.065"


.The sanitation department claims that the mean discarded plastic from all households is greater than 1.8 lb.Use a 0.05 significance to test this claim.What is the claim written in symbolic form?


1
Expert's answer
2022-05-06T12:52:41-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le 1.8"

"H_a:\mu>1.8"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=61" degrees of fredom, and the critical value for a right-tailed test is "t_c =1.67022."

The rejection region for this right-tailed test is "R = \{t: t >1.67022\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{1.911-1.8}{1.065/\sqrt{62}}\approx0.82067"

Since it is observed that "t = 0.82067< 1.67022=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for right-tailed, "df=61" degrees of freedom, "t=0.82067" is "p =0.207515," and since "p=0.207515>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is greater than 1.8, at the "\alpha = 0.05" significance level.